Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability again :(

Expert replies
by mehaksal » Mon Aug 27, 2012 7:40 am
A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?
A. 15/28 B. 1/4 C. 9/16 D. 1/32 E. 1/16

I know it's a simple question, with a simple method of solving, but with probability I tend to go very confused about which way to go.
Eg. Here I thought of going 8C2 - 2c2 way, which is very wrong!
Help me out!
Join the discussion
Source: — Problem Solving |

by cypherskull » Mon Aug 27, 2012 8:08 am
P(Not Blue) = 1 - P(Blue)

1st card

P(Blue) = 2/8 = 1/4
So, P(Not Blue) = 3/4 -------(1)

2nd card having selected one non-blue card
P(Blue) = 2/7 [Since there are still 2 blue cards]
So, P(Not Blue) = 5/7 -------(2)

Therefore, probability of selecting both blue cards = (3/4)*(5/7) = [spoiler]15/28 Ans: A[/spoiler]

Let me know the OA.
Regards,
Sunit

________________________________

Kill all my demons..And my angels might die too!
Join the discussion

by cypherskull » Mon Aug 27, 2012 8:21 am
Another easier approach could be as follows:

P(not blue) = (Number of ways a non-blue card can be selected)/(Total # of ways 2 cards can be selected
= 6C2/8C2
[spoiler]= 15/28[/spoiler]
Regards,
Sunit

________________________________

Kill all my demons..And my angels might die too!
Join the discussion